# Decomposing The Dark Matter of Sparse Autoencoders

### v1

@@ 5.3 Analyzing Per-Token Scaling Predictions
Finally, we provide a simple explanation for why per-token SAE errors are highly predictable between SAEs of different sizes. For this, we only need Eq. 9. Since $\texttt{Dense}({\mathbf{x}})$ and $\texttt{Introduced}({\mathbf{x}})$ stay mostly constant as $m$ increases, for large $m$ the SAE error stays mostly constant because it is primarily determined by these components. Thus, since $m=16k$ is already large, a linear prediction that is just a slightly smaller version of the current error performs well. Additionally, this reasoning suggests a natural experiment: if we can predict $\texttt{Introduced}({\mathbf{x}})$ on a per-token level (which we hypothesize we can do with the non-linearly predictable component of SAE error), we may be able to better predict the floor of SAE scaling and therefore better predict larger SAE errors; we run this experiment in Fig. 8, where we find an affirmative answer.
6 Analyzing Components of SAE Error
In this section, we run experiments that seek to prove our hypothesis from Section 5.2: that the split of $\texttt{SaeError}({\mathbf{x}})$ into a linearly predictable component and the non-linearly predictable component is meaningful. For convenience, given a probe ${\bm{b}}^{*}$ from Eq. 3, we write
$\texttt{LinearError}({\mathbf{x}})$ :-$\coloneq\texttt{SaeError}({\mathbf{x}})-{\bm{b}}^{*}\cdot{\mathbf{x}}:$
$\texttt{NonlinearError}({\mathbf{x}})$ :-$\coloneq\texttt{SaeError}({\mathbf{x}})-\texttt{LinearError}({\mathbf{x}}):$
Applying the Norm Prediction Test: For our first experiment, we run the norm prediction test from Eq. 2 on six different random vectors: ${\mathbf{x}}$, $\texttt{Sae}({\mathbf{x}})$, $\texttt{Sae}({\mathbf{x}})-\texttt{NonlinearError}({\mathbf{x}})$ (just the linearly predictable part of the SAE reconstruction), $\texttt{SaeError}({\mathbf{x}})$, $\texttt{LinearError}({\mathbf{x}})$, and $\texttt{NonlinearError}({\mathbf{x}})$. The results are shown as a violin plot for each component across all layer $20$ Gemma Scope SAEs in Fig. 8 (the $\texttt{Sae}({\mathbf{x}})$ bar is just a summary of Fig. 3).
Firstly, we note that $\left\lVert{\mathbf{x}}\right\rVert_{2}^{2}$ can almost be perfectly predicted from ${\mathbf{x}}$. This is reassuring news for the linear representation hypothesis, as it implies that ${\mathbf{x}}$ can indeed be well modeled as the sum of many one-dimensional features, at least from the perspective of this test.
We also find that $\texttt{NonlinearError}({\mathbf{x}})$ has a notably lower score on this test. This supports our hypothesis that unlike ${\mathbf{x}}$, this component does not consist mostly of a sparse sum of linear features from ${\mathbf{x}}$, and may be partly composed of $\texttt{Introduced}({\mathbf{x}})$.
Figure 7: Violin plot of norm prediction tests for all layer $20$ Gemma Scope SAEs. We plot the $R^{2}$ of a linear regression from ${\mathbf{x}}$ to to each random vector’s norm squared.
Figure 8: Auto-interpretability results on SAEs trained on the linear and nonlinear components of $\texttt{SaeError}({\mathbf{x}})$. “Not” represents contexts that the SAE latent did not activate on, while each $Qi$ represents activating examples from decile $i$.
Training SAEs on $\texttt{SaeError}({\mathbf{x}})$ Components: Another empirical test we run is training an SAE on $\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{LinearError}({\mathbf{x}})$. Our hypothesis is that it will be harder to learn an SAE for $\texttt{NonlinearError}({\mathbf{x}})$ than $\texttt{LinearError}({\mathbf{x}})$, since $\texttt{LinearError}({\mathbf{x}})$ should have a higher proportion of true linear features. We choose a fixed Gemma Scope layer $20$ SAE with $16k$ latents and $L_{0}\approx 60$ to generate $\texttt{SaeError}({\mathbf{x}})$ from. This SAE has nonlinear and linear components of the error approximately equal in norm and $R^{2}$ of the total $\texttt{SaeError}({\mathbf{x}})$ they explain, so it presents a fair comparison. We train SAEs to convergence (about 100M tokens) on each of these components of error and find that indeed, the SAE trained on $\texttt{NonlinearError}({\mathbf{x}})$ converges to a fraction of variance unexplained an absolute $5$ percent higher than the SAE trained on the linear component of SAE error ($\approx 0.59$ and $\approx 0.54$ respectively).
One confounding factor is that the linear component of SAE error additionally contains $\texttt{Dense}({\mathbf{x}})$, which may also be harder for the SAE to learn. Thus, we additionally examine the interpretability of the learned SAE latents using automated interpretability (this technique was first proposed by Bills et al. (2023) for interpreting neurons, and first applied to SAEs by Cunningham et al. (2023)). Specifically, we use the implementation introduced by Juang et al. (2024), where a language model (we use Llama 3.1 70b (AI@Meta, 2024)) is given top activating examples to generate an explanation, and then must use only that explanation to predict if the feature fires on a test context. Our results in Fig. 8 show that indeed, the SAE trained on linear error produces latents that are about an absolute $5\%$ more interpretable across all activation firing deciles (we average results across $1000$ random features for both SAEs, where for each feature use $7$ examples in each of the $10$ feature activation deciles as well as $50$ negative examples, and show $95\%$ confidence intervals).
Downstream Cross Entropy Loss of $\texttt{SaeError}({\mathbf{x}})$ Components: A common metric used to test SAEs is the percent of cross entropy loss recovered when the SAE reconstruction is inserted into the model in place of the original activation versus an ablation baseline (see e.g. Bloom (2024)). We modify this test to specifically examine the different components of $\texttt{SaeError}({\mathbf{x}})$: we compare the percent of the cross entropy loss recovered when replacing ${\mathbf{x}}$ with $\texttt{Sae}({\mathbf{x}})$ plus either $\texttt{LinearError}({\mathbf{x}})$ or $\texttt{NonlinearError}({\mathbf{x}})$ to the baseline of inserting just $\texttt{Sae}({\mathbf{x}})$ in place of ${\mathbf{x}}$. To estimate how much each component “should” recover, we use two metrics: the average norm of the component relative to the total norm of $\texttt{Sae}({\mathbf{x}})$ and the percent of the variance that the component recovers between $\texttt{Sae}({\mathbf{x}})$ and ${\mathbf{x}}$. The results, shown in Fig. 9, show that for the most part these metrics are reasonable predictions for both types of error. That is, both $\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{LinearError}({\mathbf{x}})$ proportionally contribute to the SAE’s increase in downstream cross entropy loss, with possibly a slightly higher contribution than expected for $\texttt{LinearError}({\mathbf{x}})$.
Using $\left\lVert\texttt{NonlinearError}({\mathbf{x}})\right\rVert$ to Predict Scaling: Following up on our discussion in Section 5.3, we are interested in whether $\texttt{NonlinearError}({\mathbf{x}})$ can help with predicting SAE per-token error norm scaling, as this might suggest that it contains a larger component of $\texttt{Introduced}({\mathbf{x}})$ than $\texttt{LinearError}({\mathbf{x}})$. Formally, we solve for
Figure 9: Results of intervening in the forward pass and replacing ${\mathbf{x}}$ with $\texttt{Sae}({\mathbf{x}})+\texttt{NonlinearError}({\mathbf{x}})$ and $\texttt{Sae}({\mathbf{x}})+\texttt{LinearError}({\mathbf{x}})$ during the forward pass. Reported in percent of cross entropy loss recovered with respect to the difference between the same intervention with $\texttt{Sae}({\mathbf{x}})$ and with the normal model forward pass.
7 Reducing $\texttt{NonlinearError}({\mathbf{x}})$
If indeed $\texttt{NonlinearError}({\mathbf{x}})$ is partly made up of $\texttt{Introduced}({\mathbf{x}})$, we may be able to reduce it by improving SAEs. Thus, in this section, we investigate to what extent simple techniques can reduce $\texttt{NonlinearError}({\mathbf{x}})$.
7.1 Using a More Powerful Encoder
Figure 10: Investigations towards reducing nonlinear SAE error.
Our first approach for reducing nonlinear error is to try improving the encoder. We use a recent approach suggested by Smith (2024a): applying a greedy inference time optimization algorithm called gradient pursuit to a frozen learned SAE decoder matrix. We implement the algorithm exactly as described by Smith (2024a) and run it on all layer $20$ Gemma Scope $9b$ SAEs closest to $L_{0}\approx 60$. For each example ${\mathbf{x}}$ with reconstruction $\texttt{Sae}({\mathbf{x}})$, we use the gradient pursuit implementation with an $L_{0}$ exactly equal to the $L_{0}$ of ${\mathbf{x}}$ in the original $\texttt{Sae}({\mathbf{x}})$.
Figure 11: Percent decrease in FVU when doing a linear prediction with the norms squared of both SAE error and SAE nonlinear error, as opposed to with just the norm squared of SAE error.
Using these new reconstructions of ${\mathbf{x}}$, we repeat Eq. 3 and do a linear transformation from ${\mathbf{x}}$ to the inference time optimized reconstructions. We then regenerate the same scaling plot as Fig. 1 and show this figure in Fig. 10(a). Our first finding is that pursuit indeed decreases the total FVU of $\texttt{Sae}({\mathbf{x}})$ by $3$ to $5\%$; as Smith (2024a) only showed an improvement on a small $1$ layer model, to the best of our knowledge we are the first to show this result on state of the art SAEs. Our most interesting finding, however, is that the $\texttt{FVU}_{\texttt{nonlinear}}$ stays almost constant when compared to the original SAE scaling in Fig. 1. In other words, if our tests are accurate, most of the reduction in FVU comes from better learning $\texttt{Dense}({\mathbf{x}})$ and reducing the linearly explainable error, not from reducing $\texttt{Introduced}({\mathbf{x}})$. In Fig. 1, we plot the additional reduction in $\texttt{FVU}_{\texttt{nonlinear}}$ as the contribution of encoder error; because $\texttt{FVU}_{\texttt{nonlinear}}$ stays almost constant, this section is very narrow.
7.2 Linear Projections Between Adjacent SAEs
Our second approach for reducing nonlinear error is to try to linearly explain it in terms of the outputs of previous SAEs. The motivation for this approach is that during circuit analysis (see e.g. Marks et al. (2024)), an SAE is trained for every component in the model, and being able to explain parts of the SAE error in terms of prior SAEs would directly decrease the magnitude of noise terms in the discovered SAE feature circuits. For the Gemma 2 architecture at the locations the SAEs are trained on, each residual activation can be decomposed in terms of prior components:
In Fig. 10(b), we plot the $R^{2}$ of a regression from each of these right hand side components to each of the different components of an SAE trained on $\texttt{Resid}_{\texttt{layer}}$ ($\texttt{SaeError}({\mathbf{x}})$, $\texttt{LinearError}({\mathbf{x}})$, and $\texttt{NonlinearError}({\mathbf{x}})$) for layer $19$-$20$ with all SAEs chosen with width $16k$ and $L_{0}\approx 60$. We find that we can explain a small amount (up to $\approx 10\%$) of total $\texttt{SaeError}({\mathbf{x}})$ using previous components, which may be immediately useful for circuit analysis.
We also find that $\texttt{SaeError}({\mathbf{x}})$ itself can explain $50\%$ of the variance in the nonlinear error, although this may not be entirely surprising, as the nonlinear error is a function of $\texttt{Sae}({\mathbf{x}})$:
These results mean that we might be able to explain some of the SAE Error using a circuits level view, but overall there are still large parts of each error component unexplained.
8 Conclusion
The fact that SAE error can be predicted and analyzed at all is surprising; thus, our findings are intriguing evidence that SAE error, and not just SAE reconstructions, are worthy of analysis. Indeed, as a byproduct of studying SAE error, we have discovered a number of interesting practical applications: we can predict which tokens will have the highest error in a large SAE without needing to train it, and we can decrease error terms in SAE circuit analysis. Additionally, the presence of constant nonlinear error at a fixed sparsity as we scale implies that scaling SAEs may not be the only (or best) way to explain more of model behavior. Future work might explore alternative penalties besides sparsity or new ways to learn better dictionaries. We note that our tests are also approximate; we argue for the existence of $\texttt{Introduced}({\mathbf{x}})$ as a separate term from $\texttt{Dense}({\mathbf{x}})$, but the exact magnitude of each component remains uncertain. Ultimately, we believe that there is still room to make SAEs better, not just bigger.
Acknowledgments
Appendix A Extra Error Prediction Experiments
A.1 Note on Feature Shrinkage
Earlier SAE variants were prone to feature shrinkage: the observation that $\texttt{Sae}({\mathbf{x}})$ systematically undershot ${\mathbf{x}}$. Current state of the art SAE variants (e.g. JumpReLU SAEs, which we examine in this work), are less vulnerable to this problem, although we still find that Gemma Scope reconstructions have about a $10$% smaller norm than ${\mathbf{x}}$. One potential concern is that the ${\mathbf{b}}^{*}$ in Eq. 3 that we learn is merely predicting this shrinkage. If this was the case, then the cosine similarity of the linear error prediction $({\bm{b}}^{*})^{T}\cdot{\mathbf{x}}$ with ${\mathbf{x}}$ would be close to $1$; however, in practice we find that it is around $0.5$, so ${\bm{b}}^{*}$ is indeed doing more than predicting shrinkage.
A.2 Breaking Apart Error Per Token
A.3 Norm Prediction Baselines
In Fig. 14, we run a linear regression from different components to $\lVert\texttt{SaeError}({\mathbf{x}})\rVert$. We find that is is not “easy” to predict SAE error norm, especially at later layers; the token identity, SAE L0, activation norm, and model loss all do significantly worse than using the full activation. It is interesting to note that at the first few layers, token identity does better at predicting SAE error than a probe of the activations; this is perhaps not surprising, since recent results from e.g. Lad et al. (2024) show that very early layers primarily operate on a per token level.
Appendix B More Info on Modeling Activations
B.1 Proof of Claim from Section 5.1
@@ Appendix B More Info on Modeling Activations
B.2 Empirical Norm Prediction
In this experiment, we aim to determine to what extent our analysis in Section 5.1 holds true in practice on almost orthogonal true SAE features. Thus, we use a random vector that we can control: $\texttt{Sae}({\mathbf{x}})$. Specifically, we learn a probe ${\bm{a}}^{*}$ for the Gemma Scope layer $20$, width $131k$, $L_{0}=62$ SAE as in Eq. 2, except with the regressor equal to $\texttt{Sae}({\mathbf{x}})$ and the target equal to $\lVert\texttt{Sae}({\mathbf{x}})\rVert:$
@@ B.2 Empirical Norm Prediction
Finally, we plot this dot product against the average latent activation in Fig. 14. If ${\bm{a}}^{*}$ indeed equals the sum of the latents weighted by their activation, as we predict in Section 5.1, then these two quantities should be approximately equal, which we indeed see in the figure.
B.3 Synthetic SAE Error Vector Experiments
@@
Figure 15: Top: When controlled amounts of noise are added to synthetic data $\texttt{Sae}({\mathbf{x}}^{\prime})$ and ${\mathbf{x}}^{\prime}$, the result is a plot similar to Fig. 1. Bottom: The nonlinear and linear error estimates (as shown at top) accurately correlate with the amount of noise added. The exact correlation between synthetic added noise and resulting estimated error components across these noise levels are shown in Table 5.2
We also tried running this test on a sparse sum of random vectors, which did not work as well, possibly due to not including the structure of the SAE vectors (Giglemiani et al., 2024); see Appendix C for more details.
Appendix C Synthetic Experiments with Random Data
For this set of experiments, we generated a random vector ${\mathbf{x}}^{\prime}$ that was the sum of a power law of $100k$ random gaussian vectors in $\mathbb{R}^{4000}$ with expected $L_{0}$ of around $100$. To simulate the SAE reconstruction and SAE error, we simply masked a portion of the vectors in the sum of ${\mathbf{x}}^{\prime}$. Unlike the more realistic synthetic data case we describe in Section B.3, this did not work as expected: even in the case with no noise added to ${\mathbf{x}}^{\prime}$ or the simulated reconstruction, the variance explained by the sum of the linear estimate of the error plus the reconstructed vectors plotted against the number of features “ablated” formed a parabola (with minimum variance explained in the middle region), as opposed to a straight line as in Fig. 15.

### v2

@@ B.3 Analyzing Per-Token Scaling Predictions
Finally, we provide a simple explanation for why per-token SAE errors are highly predictable between SAEs of different sizes. For this, we only need Eq. 20. Since $\texttt{Dense}({\mathbf{x}})$ and $\texttt{Introduced}({\mathbf{x}})$ stay mostly constant as $m$ increases, for large $m$ the SAE error stays mostly constant because it is primarily determined by these components. Thus, since $m=16k$ is already large, a linear prediction that is just a slightly smaller version of the current error performs well. Additionally, this reasoning suggests a natural experiment: if we can predict $\texttt{Introduced}({\mathbf{x}})$ on a per-token level (which we hypothesize we can do with the non-linearly predictable component of SAE error), we may be able to better predict the floor of SAE scaling and therefore better predict larger SAE errors; we run this experiment in Appendix E, where we find an affirmative answer.
Appendix C More Info on Modeling Activations
C.1 Proof of Claim from Section B.1
@@ Appendix C More Info on Modeling Activations
C.2 Empirical Norm Prediction
In this experiment, we aim to determine to what extent our analysis in Section B.1 holds true in practice on almost orthogonal true SAE features. Thus, we use a random vector that we can control: $\texttt{Sae}({\mathbf{x}})$. Specifically, we learn a probe ${\bm{a}}^{*}$ for the Gemma Scope layer $20$, width $131k$, $L_{0}=62$ SAE as in Eq. 10, except with the regressor equal to $\texttt{Sae}({\mathbf{x}})$ and the target equal to $\lVert\texttt{Sae}({\mathbf{x}})\rVert:$
@@ C.2 Empirical Norm Prediction
Finally, we plot this dot product against the average latent activation in Fig. 14. If ${\bm{a}}^{*}$ indeed equals the sum of the latents weighted by their activation, as we predict in Section B.1, then these two quantities should be approximately equal, which we indeed see in the figure.
C.3 Synthetic SAE Error Vector Experiments
@@
Figure 15: Top: When controlled amounts of noise are added to synthetic data $\texttt{Sae}({\mathbf{x}}^{\prime})$ and ${\mathbf{x}}^{\prime}$, the result is a plot similar to Fig. 1. Bottom: The nonlinear and linear error estimates (as shown at top) accurately correlate with the amount of noise added. The exact correlation between synthetic added noise and resulting estimated error components across these noise levels are shown in Table B.2
We also tried running this test on a sparse sum of random vectors, which did not work as well, possibly due to not including the structure of the SAE vectors (Giglemiani et al., 2024); see Appendix D for more details.
Appendix D Synthetic Experiments with Random Data
For this set of experiments, we generated a random vector ${\mathbf{x}}^{\prime}$ that was the sum of a power law of $100k$ random gaussian vectors in $\mathbb{R}^{4000}$ with expected $L_{0}$ of around $100$. To simulate the SAE reconstruction and SAE error, we simply masked a portion of the vectors in the sum of ${\mathbf{x}}^{\prime}$. Unlike the more realistic synthetic data case we describe in Section C.3, this did not work as expected: even in the case with no noise added to ${\mathbf{x}}^{\prime}$ or the simulated reconstruction, the variance explained by the sum of the linear estimate of the error plus the reconstructed vectors plotted against the number of features “ablated” formed a parabola (with minimum variance explained in the middle region), as opposed to a straight line as in Fig. 15.
@@ Appendix D Synthetic Experiments with Random Data
Appendix E Using $\left\lVert\texttt{NonlinearError}({\mathbf{x}})\right\rVert$ to Predict Scaling:
Following up on our discussion in Section B.3, we are interested in whether $\texttt{NonlinearError}({\mathbf{x}})$ can help with predicting SAE per-token error norm scaling, as this might suggest that it contains a larger component of $\texttt{Introduced}({\mathbf{x}})$ than $\texttt{LinearError}({\mathbf{x}})$. Formally, we solve for
Figure 16: Percent decrease in FVU when additionally using the squared norms of nonlinear error to predict SAE error norm.